What is the length and width of a rectangle that is 80m? The length is 7m more than twice the width. Find the dimensions.
Answer by london maths tutor(243)

Let 's the length of the rectangle be "L" and the width of the rectangle be "w".
L = 7 + 2w (The length is 7m more than twice the width.)
Perimeter = 80 m
Perimeter = 2L + 2w = 80 m
2(7+2w) + 2w = 80m
14 + 4w + 2w = 80m
6w = 80 - 14 m
6w = 66 m
w = 66/6 m
w = 11 m
Put w = 11 into L = 7 + 2w
L = 7 + 2(11) m
L = 7 + 22 m
L = 29 m
Answer: Length = 29m and width = 11m